Find all the zeros of the quadratic polynomial f(x)=abx^2+(b^2-ac)x-bc

Grade:

2 Answers

pankaj kumar raman
11 years ago

abx²+(b²-ac)x-bc=0

x=-(b²-ac)±√[(b²-ac)²+4.ab²c]  ÷ 2ab ..................(1)

x=-(b²-ac)±(b²+ac)  ÷  2ab

x=-b²/ab , ac/ab

x= -b/a , c/a

 

 

eq (1) can be solved like this

√[bΛ4 + a²c² - 2ab²c + 4ab²c] = √[bΛ4 + a²c² + 2ab²c] = √[b²+ac]² = b² + ac

Vedant Vedant
11 years ago

Thanks!!